### v1

# Learning Logical Rules using Minimum Message Length

Abstract
Unifying probabilistic and logical learning is a key challenge in AI. We introduce a Bayesian inductive logic programming approach that learns minimum message length programs from noisy data. Our approach balances hypothesis complexity and data fit through priors, which explicitly favour more general programs, and a likelihood that favours accurate programs. Our experiments on several domains, including game playing and drug design, show that our method significantly outperforms previous methods, notably those that learn minimum description length programs. Our results also show that our approach is data-efficient and insensitive to example balance, including the ability to learn from exclusively positive examples.
Introduction
Inductive logic programming (ILP) is a form of machine learning. The goal is to induce a logic program (a set of logical rules) that generalises examples and background knowledge (BK) (Muggleton 1991; Cropper and Dumancic 2022).
As with all forms of machine learning, ILP learns from noisy data. A popular approach is to search for a minimal description length (MDL) (Rissanen 1978, 1998, 1987; Grünwald 2007) hypothesis (Quinlan 1990; Muggleton 1995; Srinivasan 2001; Cabalar, Pearce, and Valverde 2007; Hocquette et al. 2024). MDL is a mathematical formalisation of Occam’s razor: a preference for simpler (formalised as shorter) explanations of data. The goal of MDL is to encode data as a binary string as concisely as possible.
Minimum message length (MML) (Wallace 2005) is another formalisation of Occam’s razor. The goal of MML is to find the most concise explanation of data, where an explanation is a two-part message consisting of a hypothesis (the first part) followed by the data given that the hypothesis is true (the second part).
MDL and MML are closely related, but differ in philosophy, mathematical grounding, and practical usage (Baxter and Oliver 1994; Makalic and Schmidt 2022). We briefly summarise the key differences, but expand on them in Section 2. MML is explicitly Bayesian and uses priors over hypotheses and integrates them into encoding length. MDL avoids priors and often treats hypotheses as tools for encoding data rather than probabilistic entities. In MML, the message is two-part, first asserting the hypothesis, followed by data given that the asserted hypothesis is true. In MDL, the message may be one-part and often does not encode the hypothesis itself. MML and MDL also handle continuous parameters differently (Baxter and Oliver 1994; Grünwald 2007).
In this paper, we use the MML principle to learn logic programs from noisy examples. Specifically, we design an encoding scheme which allows for lossless communication of a probabilistic logic program (De Raedt, Kimmig, and Toivonen 2007) and examples. The cost of a hypothesis is the length of the hypothetical message which would result if we were to communicate the examples using this hypothesis.
Our main claim is that MML outperforms a previous MDL-based score (Hocquette et al. 2024), shown to achieve state-of-the-art performance when learning logic programs from noisy examples. We term this previous score C-MDL. We argue that MML is better suited for this task for two reasons: (i) it uses prior probabilities, allowing us to favour more general programs, and (ii) it handles continuous parameters in a way that provides automatic regularisation (Wallace 2005; Farr and Wallace 2002; Wallace and Freeman 1987). MML has two more advantages over C-MDL: (iii) its hypothesis cost corresponds to an exact message length, and (iv) it uses a structured likelihood function that penalises over-generalisation, with this penalty being regulated by the generality priors from (i). Whilst an exact MDL approach could incorporate (iii) and (iv), it would still lack the intrinsic MML benefits of (i) and (ii).
Specifically, we claim:
Claim 1:
MML performs similarly to C-MDL on general tasks
Claim 2:
MML outperforms C-MDL when learning from different example proportions, including learning from exclusively one type of example
Claim 3:
MML requires less data than C-MDL
Claim 4:
An explicit prior preference for more general programs improves the data efficiency of MML
Claim 5:
MML performs similarly to C-MDL on noisy data
Claim 6:
A practical approximation to MML outperforms C-MDL when learning from unbalanced examples
Contributions
Overall, our contributions are:
  1. 1.
    We formulate the program learning problem in Bayesian and MML terms. We use the MML framework to derive a cost function for learning probabilistic logic programs, consistent with the Problog semantics (De Raedt, Kimmig, and Toivonen 2007).
  2. 2.
    We introduce a practical method to learn approximate MML programs from examples.
  3. 3.
    We empirically show that the MML score outperforms C-MDL on predictive accuracy and matches its performance on noisy data. The method excels when learning from few or unbalanced examples.
Related Work
ILP. Most ILP approaches learn from noisy data (Quinlan 1990; Muggleton 1995; Blockeel and De Raedt 1998; Srinivasan 2001). For instance, MaxSynth (Hocquette et al. 2024) searches for programs that minimise the total of the training errors and literals in the program. In contrast to our approach, these approaches do not allow for the explicit declaration of prior beliefs regarding the quantity of noise in the examples, do not employ techniques to mitigate the effect of example balance in their learning, and do not return probabilistic logic programs.
Positive only learning. While several ILP approaches have been proposed that learn from exclusively positive examples (De Raedt and Bruynooghe 1993; Quinlan and Cameron-Jones 1995; Mooney and Califf 1995), most require an assumption of example completeness. One exception (Muggleton 1996) also models the program learning task under a Bayesian framework, but is restricted to noiseless data, uses coarse size-related priors, and employs a greedy search to find the hypothesis of maximum posterior probability. We differ by handling noise and using a prior that exactly corresponds to the program message length.
MML and MDL. MML and MDL have been applied to learn a wide range of models, including neural networks (Makalic, Allison, and Dowe 2003; Hinton and Van Camp 1993), decision trees (Wallace and Patrick 1993), automata (Wallace 2005), and statistical distributions (Farr and Wallace 2002). The most relevant prior work is (Conklin and Witten 1994), which learned deterministic logic programs. However, their coding scheme is inefficient and cannot handle noisy data. We introduce a more principled coding scheme, an improved search procedure, and extend the framework to learn probabilistic logic programs. We choose MML over MDL for its Bayesian framework and method of handling parameter uncertainty: encoding real-valued parameters to a finite precision. MDL usually uses approximations or penalised maximum likelihood (ML) estimators to estimate continuous parameters. MML’s handling often yields less biased and consistent estimators that are often more accurate than ML (Baxter and Oliver 1994; Wallace 2005).
Probabilistic logic programming. Previous work has focused on learning rule probabilities (Gutmann, Thon, and De Raedt 2011) or learning rules from Bayesian networks (De Raedt and Kersting 2008). Probfoil (De Raedt et al. 2015) is built on the Problog semantics (De Raedt, Kimmig, and Toivonen 2007) and learns probabilistic logic programs, but is frequentist, employs a greedy search, and requires both positive and negative examples. Markov logic networks (MLN) (Richardson and Domingos 2006) are another approach to unify probability and logic. MLNs are first-order knowledge bases with a weight attached to each logical formula and consider the joint distribution of a set of predicates. Our approach is Bayesian, uses a constraint solver in the search, can learn from exclusively positive examples, learns a single hypothesis, and employs MML as the hypothesis selection criterion.
Problem Setting
Inductive Logic Programming
We formulate our approach in the ILP learning from entailment setting (De Raedt 2008). We define an ILP input:
Definition 1 (ILP input).
An ILP input is a tuple $(E,B,\mathcal{H})$ where $E=(E^{+},E^{-})$ is a pair of sets of ground atoms denoting positive ($E^{+}$) and negative ($E^{-}$) examples, $B$ is background knowledge, and $\mathcal{H}$ is a hypothesis space, i.e a set of possible hypotheses.
We restrict the hypothesis space and background knowledge to definite programs with the least Hebrand model semantics (Lloyd 2012). We define a cost function:
Definition 2 (Cost function).
Given an ILP input $(E,B,\mathcal{H})$, a cost function $cost_{E,B}~:$$\mathcal{H}~\mapsto$$\mathbb{N}$ assigns a numerical cost to each hypothesis in $\mathcal{H}$.
Given an ILP input and a cost function $cost_{E,B}$, we define an optimal hypothesis:
Definition 3 (Optimal hypothesis).
Given an ILP input $(E,B,\mathcal{H})$ and a cost function costE,B, a hypothesis $H\in\mathcal{H}$ is optimal with respect to costE,B when $\forall h^{\prime}\in\mathcal{H}$, costE,B($H$) $\leq$ costE,B($H^{\prime}$).
We describe the C-MDL cost function used by Hocquette et al. (2024):
Definition 4 (C-MDL cost function).
Given a hypothesis $H$ containing $size(H)$ literals which results in $fp(H)$ false positives and $fn(H)$ false negatives on training examples, the C-MDL cost of this hypothesis is: $cost_{C\text{-}MDL}(H)=size(H)+fp(H)+fn(H)$.
Next, we introduce an MML-based cost function.
Minimum Message Length
MML aims to find the most concise two-part explanation of examples $E$. An explanation consists of a hypothesis $H$ (first part) and examples given the hypothesis is true $E|H$ (second part). We briefly describe this communication format, following Wallace and Patrick (1993).
A hypothesis $H$ can optimally encode examples $E$ as a binary string of length:
$C(E|H)=-\log_{2}{P(E|H)}$ (1)
Where $P(E|H)$ is the probability of $E$ given that $H$ is true. This string is decodable to a receiver with knowledge of $H$. If $H$ is unknown, it must be communicated to the receiver. One way is to prepend the example string with a second binary string that communicates $H$ in a code agreed a priori. The length of this string is $C(H)=-\log_{2}{P(H)}$. The total message length is:
$C(H,E)=C(H)+C(E|H)$ (2)
Equation 2 is the MML length function, which balances hypothesis complexity and data fit. A complex hypothesis must explain the data succinctly to be selected over a simple hypothesis that less succinctly explains the data. An MML optimal hypothesis is one that has the smallest value of $C(H,E)$.
We now describe how to define the message lengths for each part of Equation 2.
Hypothesis Message Length: $C(H)$
We now describe the type of hypothesis we communicate.
Probabilistic Logic Programs
We learn a set of rules $R$ that generalise the relation p/a from training examples. We use these rules to predict the probability that an example is true or false. We describe each case with a separate probabilistic rule with Problog semantics (De Raedt, Kimmig, and Toivonen 2007):
Case 1: When $R$ entails an example, the example is true with probability $\theta_{+}$ and false with probability $1-\theta_{+}$, described with the probabilistic rule:
$\theta_{+}::\texttt{predTrue(}X_{1},\dots,X_{a}\texttt{):-p(}X_{1},...,X_{a}\texttt{).}$
An example is predicted to be true when the probability of $\texttt{predTrue(}X_{1},\dots,X_{a}\texttt{)}$ is greater than $0.5$.
Case 2: When $R$ does not entail an example, the example is false with probability $\theta_{-}$ and true with probability $1-\theta_{-}$, described with the probabilistic rule:
$\theta_{-}::\texttt{predFalse(}X_{1},\dots,X_{a}\texttt{):-not }\texttt{p(}X_{1},\dots,X_{a}\texttt{).}$
We need two probabilistic rules because cases 1 and 2 are mutually exclusive, so they must be described separately. As these cases also cover all example possibilities, the union of their respective rules provides a probabilistic prediction for any example. We take the union of these rules with $R$ as our MML hypothesis $H$. We denote the probabilistic rules corresponding to cases 1 and 2 of the hypothesis $H$ as $H_{+}$ and $H_{-}$, respectively. We only send $R$, $\theta_{+}$, and $\theta_{-}$ in the message and assume the receiver knows how to construct $H$ from this information.
Example Group Size Length:
Each sub-hypothesis $H_{+}$ and $H_{-}$ communicates the examples that they respectively entail. Different sub-hypotheses entail different training examples, so partition the example set into two disjoint subsets. $H_{+}$ communicates the true positives and false positives. $H_{-}$ communicates the true negatives and false negatives. These subsets are sent separately, so their sizes must be communicated to ensure that the receiver knows how many examples to expect per set. We define $tp(H)$, $fp(H)$, $tn(H)$, $fn(H)$ as the total number of true positives, false positives, true negatives and false negatives predicted by the hypothesis $H$ on the training examples, respectively. Let $N$ be the total number of training examples known a priori to the receiver. The sender sends the total number of positive predictions $k=tp(H)+fp(H)$. The receiver infers the number of negative predictions $tn(H)+fn(H)$ as NkN-k. The length of communicating a particular $k$ value is based on its probability, which we model as the outcome of two independent binomial trials. These trials have a known prior probability of classification error $r$ and result in hypothetical classifications of the training examples into $tp$, fpfp, tntn, fnfn true positives, false positives, true negatives and false negatives, respectively.
  1. 1.
    Trial 1: classifying the $tp+fn$, positive examples. A trial of $tp+fn$ flips, where each success corresponds to a $tp$ and each failure, an $fn$.
  2. 2.
    Trial 2: classifying the $tn+fp$, negative examples. A trial of $tn+fp$ flips, where each success corresponds to a $tn$ and each failure, an fpfp
Then the probability of a specific, hypothetical $tp,fp$ outcome is the product of the probabilities of these two trials:
$P(tp,fp)$ $=\binom{tp+fn}{tp}(1-r)^{tp}r^{fn}$ (3)
$\times\binom{tn+fp}{tn}(1-r)^{tn}r^{fp}$
The probability of $k$ positive predictions is the sum of the probabilities of all $(tp,fp)$ outcomes such that $tp+fp=k$. The length of sending $k$ is the negative log of this probability.
Hypothesis Syntax Length: $C_{R}(H)$
The receiver can deduce the complete hypothesis $H$ from (i) the set of rules $R$ that generalise the examples and (ii) the probabilities with which each subhypothesis holds ($\theta_{+}$ and $\theta_{-}$). We define three functions which map a hypothesis $H$ to these components: $\mathbf{R}(H)=R$, $\mathbf{\Theta_{+}}(H)=\theta_{+}$ and $\mathbf{\Theta_{-}}(H)=\theta_{-}$. We first describe the communication of the syntax of $R$, which we encode piecewise. We use the example of communicating the syntax of the rule:
         {p(X,Y):-p1(X,Y), p2(X).}
We assume that all variables in the head literal are distinct and that the head literal is already known to the receiver from the task. Therefore, only body literals need to be communicated. Communication of $R$ begins by stating the total number of body literals $l(R)$ in the set of rules, where $l(R)\in\{0,1,\dots,M\}$ and $M$ is the maximum allowed number of literals. We assume no literal number is more likely than any other, so we use a uniform code over the possible $l(R)$ values and thus ignore it in the message length. After specifying the total number of body literals $l(R)$, the sender communicates how they are grouped into rules. The number of possible literal groupings is given by the integer partition function $partition(l(R))$. In our example, there are two possible groupings:
    1) p(X,Y):-?,?.    2) p(X,Y):-?.
                          p(X,Y):-?.
This part of the message has length $\log{partition(l)}$ and is larger for programs with more literals.
After sending the literal groupings, the sender sends each rule $r\in R$ separately. The sender first communicates the predicate symbols of the literals in the body of $r$. To communicate the predicate symbols, we model their probability. The receiver knows a priori the probability of seeing a predicate symbol. We calculate these probabilities using the Herbrand Base, which is the set of all literals that can be constructed from the examples and background knowledge in a given task that is consistent with the given syntactic bias (Nienhuys-Cheng and de Wolf 1997). Let $o_{p}$ be the number of atoms that contain the predicate symbol $p$ in a Herbrand Base $HB$. Then the probability of the predicate symbol $p$ is $P_{pred}(p)=\frac{o_{p}}{|HB|}$. Therefore, more general predicate symbols have greater probabilities. We only consider non-recursive programs, so we omit the counts of head predicates in the computation of $P_{pred}(p)$. Let $ps(r)$ be the predicate symbols in the body of the rule $r$ and $fr(p,r)$ be the number of times the predicate symbol $p$ appears in the body of $r$. Then the probability of seeing the predicate symbols in $r$ is the multinomial formula:
$P_{preds}(r)=\frac{\left[\sum_{p\in ps(r)}\text{fr}(p,r)\right]!}{\prod_{p\in ps(r)}\text{fr}(p,r)!}\prod_{p\in ps(r)}\!\!\!P_{\text{pred}}(p)^{\text{fr}(p,r)}$ (4)
The message length required to communicate the predicate symbols in a rule is $\log{P_{preds}(r)}$.
After sending the predicate symbols, the sender sends the body variables. The sender first sends the frequency of each distinct variable in the rule. Let $\text{fr}(v_{i},r)$ be the number of times the variable $v_{i}$ appears in rule $r$. We define var_slots_body(r)var\_slots\_body(r) as the total arity of all the body literals in rule $r$ and d_vars(r)d\_vars(r) as the number of distinct variables in $r$. Then the number of valid $\text{fr}(v_{i},r)$ assignments to each of the d_vars(r)d\_vars(r) distinct variables in the rule, n_freqs(r)n\_freqs(r), is given by stars and bars (Ross 1976):
$n\_freqs(r)=\left(\begin{smallmatrix}d\_vars(r)+var\_slots\_body(r)-1\\ d\_vars(r)-1\end{smallmatrix}\right)$ (5)
The sender specifies the variable frequencies with a message of length $\log{n\_freqs(r)}$.
To complete the rule, the sender communicates the specific sequence of variables that fill the empty body variable slots. The permutations of identical variables or relabelling of body-only variables (such as Y and Z below) produce semantically equivalent programs.
 h(X):-p1(Y),p2(Z). = h(X):-p1(Z),p2(Y).
Accounting for equivalence, the number of possible variable sequences n_seqs(r)n\_seqs(r) in a rule with d_body_vars(r)d\_body\_vars(r) distinct body-only variables, where variable $j$ is repeated $\text{fr}(v_{j},r)$ times is:
$n\_seqs(r)=\frac{var\_slots\_body(r)!}{d\_body\_vars(r)!\prod_{i}^{d\_vars(r)}\text{fr}(v_{i},r)!}$
The sender specifies a single variable sequence with a message of length $\log{n\_seqs(r)}$. Then the message length required to communicate $R$ is:
$C_{R}(H)=\log{partition(l(\mathbf{R}(H)))}+C_{R^{\prime}}(H)$ (6)
$C_{R^{\prime}}(H)=\sum_{r\in\mathbf{R}(H)}\Big{[}$- $\log{P_{preds}(r)}+\log{n\_seqs(r)}$ (7)
$+\log{n\_freqs(r)}\Big{]}$
It is convenient to delineate $C_{R^{\prime}}(H)$ as it has additive linearity properties, which we later use in our optimisation.
Theta Length: $C_{\theta_{+},\theta_{-}}(H)$
We describe communication of the $\theta_{+}=\mathbf{\Theta}_{+}(H)$ and $\theta_{-}=\mathbf{\Theta}_{-}(H)$ values, the Bernoulli probabilities with which $H_{+}$ and $H_{-}$ each hold. An MML approximation (Wallace and Freeman 1987) to the message length of communicating these values is known. We state this length and refer the reader to the full derivation in Wallace (2005). Let $\alpha$ and $\beta$ represent the parameters of the beta prior distribution $\mathcal{B}(\alpha,\beta)$ over each $\theta$ parameter. Then the length of communicating $\theta_{+}$ is $C_{\theta_{+}}(H)$:
$(\alpha-\frac{1}{2})\log{\mathbf{\Theta}_{+}(H)}-(\beta-\frac{1}{2})\log{\big{[}1-\mathbf{\Theta}_{+}(H)\big{]}}+K$ (8)
Where $K\approx-\log{\mathcal{B}(\alpha,\beta)}-0.7425+\log{\binom{tp(H)+fp(H)}{tp(H)}}$, and:
$\mathbf{\Theta}_{+}(H)=\frac{tp(H)+\alpha-0.5}{tp(H)+fp(H)+\alpha+\beta-1}$ (9)
Similarly, the length of communicating $\theta_{-}$ is $C_{\theta_{-}}(H)$:
$(\alpha-\frac{1}{2})\log{\mathbf{\Theta}_{-}(H)}-(\beta-\frac{1}{2})\log{\big{[}1-\mathbf{\Theta}_{-}(H)\big{]}}+K$ (10)
Where $K\approx-\log{\mathcal{B}(\alpha,\beta)}-0.7425+\log{\binom{tn(H)+fn(H)}{tn(H)}}$, and:
$\mathbf{\Theta}_{-}(H)=\frac{tn(H)+\alpha-0.5}{tn(H)+fn(H)+\alpha+\beta-1}$ (11)
The message length required to communicate the theta values is the sum of the message lengths required to communicate $\theta_{+}$ and $\theta_{-}$:
$C_{\theta_{\pm}}(H)=C_{\theta_{+}}(H)+C_{\theta_{-}}(H)$ (12)
Total Hypothesis Length
The total message length of the hypothesis $C(H)$ is the sum of equations 6 and 12, and the length of communicating the sizes of the example sets:
[…]
Where $k$ is the size of the group communicated by $H_{+}$.
Examples Length: $C(E|H)$
We describe how to send examples $E$ using the hypothesis $H$. We define a likelihood function $P(E|H)$ which determines $C(E|H)$ (Equation 1). Each example is composed of a literal and a truth value. We describe sending each in turn.
Literals
We assume that no particular example is a priori more likely to be seen than any other. Let $e(H_{i})$ be all literals in the Herbrand Base entailed by the subhypothesis $H_{i}$, $i\in\{+,-\}$. Then the probability of observing a particular set of examples is the chance of randomly selecting $tp(H)+fp(H)$ examples from $e(H_{+})$ and $tn(H)+fn(H)$ examples from $e(H_{-})$:
$P_{lits}(E|H)=\left[\binom{|e(H_{+})|}{tp(H)\!+\!fp(H)}\binom{|e(H_{-})|}{tn(H)\!+\!fn(H)}\right]^{-1}$
The length of this message is $\log{P_{lits}(E|H)}$.
Truth Values
According to Problog semantics, the probability of a sequence of $tp(H)$, $fp(H)$, $tn(H)$, and $fn(H)$ examples is given by the product of two Bernoulli distributions:
$P_{T/F}(E|H)=\theta_{+}^{tp(H)}(1-\theta_{+})^{fp(H)}\times\theta_{-}^{tn(H)}(1-\theta_{-})^{fn(H)}$
The length of this message is $\log{P_{T/F}(E|H)}$
Total
The length of the message required to communicate the examples is therefore:
$C(E|H)=-\log{P_{lits}(E|H)}-\log{P_{T/F}(E|H)}$ (14)
Total Message Length
The receiver now has all the necessary information to reconstruct both the communicated hypothesis and data. The full message length is the sum of Equations 13 and 14. We search for the hypothesis $H$ which minimises $C(H,E)$, our cost function.
Algorithms
To test claims 1-6, we need algorithms that (i) compare exact MML and MDL programs and (ii) efficiently learn programs with good MML and MDL scores. Our contribution in this section is two such algorithms: RandomLearner and ApproxLearner.
RandomLearner
As the exact MML score is difficult to optimise, this learner calculates exact MML and MDL scores for a large number of random programs, selecting programs with the smallest MML and MDL scores. We generate $T$ random programs for each task. The method of generating the random programs follows. For each size 1Max_Size1-Max\_Size, we generate a maximum of $I$ rules of this size which conform to the syntactic bias given the task. Rules are constrained to only contain at most v_maxv\_max variables. From these rules, we generate a random program. We choose a random number $c$ between 1 and Max_ClausesMax\_Clauses, and randomly select $c$ rules from the set of all generated rules. This random set of rules is a random program. We repeat this $T$ times to generate all random programs. Specific values for T,Max_Size,I,v_max,c,T,\ Max\_Size,\ I,\ v\_max,\ c, and Max_ClausesMax\_Clauses may be found in Section 6.
ApproxLearner
This learner searches for programs which minimise a user-specified cost function. In this paper, this cost is either C-MDL or an approximation of MML. The ApproxLearner learning procedure is similar to the inductive ILP system MaxSynth, which minimises the C-MDL cost function (Hocquette et al. 2024). Both procedures (i) generate individual rules, (ii) test and cache the size and performance statistics of each rule on training data, and (iii) find good combinations of these rules (programs) according to the cost. We describe these stages next.
Constraints: The main difference between ApproxLearner and MaxSynth is step (i). When generating rules, MaxSynth uses C-MDL-specific constraints to efficiently prune rules from the rule space that cannot be in a C-MDL optimal hypothesis. This pruning vastly shrinks the search space and enables MaxSynth to learn globally optimal programs efficiently. Conversely, ApproxLearner uses weaker, cost function independent constraints (pruning generalisations of inconsistent rules) and is much less efficient. In the absence of MML-specific constraints, we use the ApproxLearner procedure to compare MML and C-MDL costs fairly, leaving MML-specific constraints to future work.
Search: To implement the search for good rule combinations in step (iii), both MaxSynth and ApproxLearner use constraint solvers, specialised tools which efficiently search large combinatorial spaces given a cost to optimise. ApproxLearner uses the CP-SAT (Perron and Didier 2025) constraint solver. CP-SAT requires a linear cost to perform an effective search. The C-MDL cost is already linear and is modelled directly in CP-SAT. The MML cost is not linear, so we use a linear approximation.
Linear MML Approximation: We cache the values of the length of the rule-specific information $C_{R^{\prime}}(H)$ (equation 7) for each rule, as it behaves linearly: $C_{R^{\prime}}(H_{1}\cup H_{2})=C_{R^{\prime}}(H_{1})+C_{R^{\prime}}(H_{2})$ for any two hypotheses $H_{1},H_{2}$. For all other terms in the $C(H,E)$ sum (equation 2), we select five breakpoints and linearly interpolate between each to determine a piecewise linear (PL) approximation. We provide the PL approximations, the cached $C(R)$ values, and rule statistics to the CP-SAT solver so that it may implement the search procedure.
Herbrand Base Calculations
For efficiency, in the RandomLearner experiments, the Herbrand Base is calculated once from the original trial and reused for any trial variation (altering training example size, noise level, or example balance). In the ApproxLearner experiments (Q6 below), however, the Herbrand Base is recalculated for all trials. The Herbrand Base is used to calculate the generality predicate priors and the set of possible example literals, which are all literals in the Herbrand Base that contain the example relation.
Experiments
We now test our claims. Unless otherwise mentioned, all experiments use RandomLearner. We first try to answer the question:
  1. Q1
    Does MML perform similarly to C-MDL on general tasks?
We answer Q1 by comparing the testing accuracy of MML and C-MDL programs on 30 diverse datasets.
To test our claim that our MML approach can learn from unbalanced examples, we try to answer the question:
  1. Q2
    Does MML outperform C-MDL when learning from different example proportions?
We answer Q2 by comparing the testing accuracy of MML and C-MDL programs on datasets with different proportions of examples, ranging from 10:0 (100% positive examples), 8:2, 6:4, …, 0:10 (100% negative examples). We randomly generate these datasets from those used in Q1. We restrict dataset sizes to 20 and 50 to control for the effect of the number of training examples.
To test our claim that our MML approach is more data efficient than C-MDL, we try to answer the question:
  1. Q3
    Can MML learn from fewer training examples than C-MDL?
We answer Q3 by comparing the testing accuracy of MML and C-MDL programs on training datasets of eight different training example sizes: 1, 5, 10, 20, 50, 100, 200, and 500. We build these datasets by random sampling from the original trial dataset.
To test our claim that the explicit preference for more general programs improves the data efficiency of our MML approach, we try to answer the question:
  1. Q4
    Do generality priors allow MML to learn from fewer training examples?
We answer Q4 by comparing the testing accuracy of our standard MML procedure (with generality predicate priors) against MML with uniform predicate priors on the same datasets as Q3.
To test our claim that our MML method performs similarly to C-MDL on noisy data, we try to answer the question:
  1. Q5
    Can MML handle similar amounts of noise as C-MDL?
We answer Q5 by introducing different amounts (10%, 20%, … 50%) of random noise into each of the 30 datasets and comparing the testing accuracies of the MML and C-MDL selected programs. We also run MML vs C-MDL on the real-world noisy alzheimer datasets. We introduce noise by randomly selecting a subset of training examples, which we randomly reassign as positive or negative.
To test our claim that an approximation to the MML cost function learns from unbalanced examples, we try to answer the question:
  1. Q6
    Can a piecewise approximation to the MML cost enable learning from unbalanced examples?
We answer Q6 by comparing C-MDL and MML on the noiseless datasets in Q1 and on modifications of these datasets containing only the positive or negative examples using ApproxLearner.
Experimental Choices
We make the following experimental choices.
Theta Priors We use identical beta prior distributions with $\alpha$ and $\beta$ parameters for our $\theta_{+}$ and $\theta_{-}$ values chosen to ensure that the learned $\theta$ values are always greater than $0.5$. This constraint has the following effect. When $\theta_{+}>0.5$, if our learned program holds, the relation that we describe is also likely to hold. When $\theta_{-}>0.5$, if our learned program does not hold, the relation that we describe is also likely not to hold. Constraining $\theta_{+},\theta_{-}>0.5$ results in the standard, non-negated program interpretation. The expected $\theta$ value is $e=\frac{\alpha}{\alpha+\beta}$. The probability of error $r$, used in $C(H_{+/-})$ calculation (Equation 3) is r=1er=1-e. For noiseless tasks, we assume $r$ is near zero: $\alpha=1000,000$ and $\beta=1$. For noisy tasks, we assume weaker, but still discriminatory programs: $\alpha=5000$, $\beta=1$ . When $\beta=1$, choosing $\alpha>N+1$, where $N$ is the number of training examples, enforces our constraint, which is satisfied for our largest considered training dataset (<2000<2000 examples).
Figure 1: Balanced MML and C-MDL program accuracy differences when changing the proportion of positive examples for training datasets of size 20. Blanks indicate proportions not possible with 20 examples. Error bars represent standard error.
RandomLearner Parameters We generate $10,000$ random programs for each task. Each rule has a maximum rule size of $4$ and at most $6$ variables. For each size, the maximum number of rules generated is $10,000$. Each random program contains at most $5$ rules.
ApproxLearner Parameters We allow each CP-SAT call 180 seconds to find a good rule combination. We use a timeout of 1000 seconds for each trial. The best program found before timeout is returned. The score of the empty program provides the starting point for the search.
Setup We run all RandomLearner experiments on an M2 Macbook Air (16GB RAM, 8 core CPU) and ApproxLearner programs on a server (36 core CPU, 125GB RAM). We repeat all experiments 10 times and calculate the mean and standard error. We measure balanced testing accuracy for each task to account for imbalanced datasets. All significance claims are determined by Benjamini-Hochberg corrected Wilcoxon signed-rank tests at an overall significance level of 0.05. We make all comparisons between tasks, not trials within a task, to satisfy the test’s independence assumptions.
Datasets
We test our claims on the following datasets:
Alzheimer. Real-world noisy datasets, learning attributes of drug molecules for alzheimers treatment (King et al. 1992).
Zendo. An inductive game where the goal is to determine an underlying true rule (Cropper and Hocquette 2023).
1D-ARC. A one-dimensional adaptation (Xu et al. 2023) of the ARC-AGI dataset (Chollet 2019).
Trains. Synthetic datasets whose goal is to determine a rule which differentiates eastbound from westbound trains (Larson and Michalski 1977).
Results
The experimental results follow.
Q1: Overall Performance: RandomLearner
On the 30 standard tasks, we find no statistically significant difference between MML and C-MDL. We hypothesise that this is because most datasets are balanced and/or moderately sized, qualities which allow C-MDL and standard ILP systems to perform well. These results support Claim 1.
Q2: Example Balance: RandomLearner
For training sets of size 20 and 50, the mean pairwise balanced accuracy difference consistently favours MML across all 12 tested size-example proportion datasets. This advantage is statistically significant ($p<0.05$) in 11/12 conditions. Figure 1 illustrates this trend for the size 20 datasets. One reason why MML outperforms C-MDL is because C-MDL struggles when learning from 100% positive examples. In this setup, a program that classifies all examples as true is C-MDL optimal (it makes no errors and is minimal in size). By contrast, the MML likelihood (Equation 1) penalises overly general programs. Therefore, a program which entails all possible examples results in the cost of Equation Literals as $\log\binom{|T|}{tp+fp}$ which, when $|T|>>tp+fp$, is very large. MML prefers a program which entails only the seen examples (and makes no/few errors), as this cost is instead $\log\binom{tp+fp}{tp+fp}=0$. These results support Claim 2.
Q3: Data Efficiency: RandomLearner
MML achieves a statistically significant improvement in the mean pairwise accuracy difference over C-MDL at all eight tested training sizes. On 5 out of the 8 tested sizes, MML results in more than a 4% average improvement over all tasks. The improvement is largest for datasets of size 1, yielding an improvement of $6.8\pm 2.4\%$. We hypothesise this improvement is due to our generality prior coupled with the more informative likelihood. These results support Claim 3.
Q4: Generality Priors: RandomLearner
The standard MML procedure with generality priors achieves a statistically significant ($p<0.05$) improvement of $6.6\pm 2.6\%$ in mean pairwise accuracy difference over MML with uniform priors when learning from a single example. We find no significant difference at any larger training size. These results support Claim 4.
Q5: Noise Tolerance: RandomLearner
On datasets with 10-50% added noise or the standard alzheimers datasets, we find no statistically significant difference between MML and C-MDL. We hypothesise that MML’s performance in these low-signal domains is hindered by a prior that is overly optimistic, and would be improved by one which is more carefully selected. These results support Claim 5.
Q6: Example Balance Tolerance and Overall Performance: ApproxLearner
On datasets consisting of exclusively positive and exclusively negative examples, the approximate MML cost significantly ($p<0.01$) outperforms C-MDL by $9.2\pm 2.7$% and $25.9\pm 4.0\%$ respectively. The maximum improvement was $50\%$ in both experiments: the maximum possible. The maximum losses were much smaller, $3.6\%$ and $0.0\%$ for $100\%$ and $0\%$ positive examples, respectively. Experiments on the standard noiseless tasks reveal no significant difference between MML and C-MDL. This result is despite the difficulty of optimising the more complex MML cost function. These results support Claim 6
Conclusions and Limitations
We derived an ILP cost function that is data-efficient, insensitive to example imbalance, and noise-tolerant. Our cost function enables learning from exclusively positive examples, a common failure mode for many ILP systems. Empirically, our cost outperforms the score from the state-of-the-art MDL system MaxSynth on data efficiency and example imbalance, and matches its performance on noisy data. Our introduced prior preference for more general predicates results in significantly increased data efficiency relative to a uniform prior when learning from a single example. Unlike MaxSynth, our system does not guarantee optimal programs. Future work should identify tight constraints to enable efficient optimal search to enable equivalently powerful MML-based systems. The described approach has only been demonstrated for finite domains, but may be extended to infinite domains by constructing a truncated Herbrand Base from only the ground terms observed. Alternative priors should also be explored, such as extending the presented predicate generality priors to incorporate domain knowledge. We acknowledge that the beta prior also requires user specification, which is not needed for MDL. One should instead determine robust noisy default beta priors by grid search over many datasets. This should improve performance, particularly in high-noise domains.
Acknowledgements
Ruben Sharma was supported by the EPSRC Centre of Doctoral Training in Automated Chemical Synthesis Enabled by Digital Molecular Technologies (Syntech, EP/S005722/1). Sebastijan Dumančić was supported by NWO (OCENW.M.21.325 and NWA.1418.22.008). Ross D. King was supported by the EPSRC (EP/R022925/2, EP/W004801/1 and EP/X032418/1), and by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Alice Wallenberg Foundation. Andrew Cropper was supported by his EPSRC fellowship (EP/V040340/1).